OpenStudy (anonymous):

Solve 6 and one-ninth minus 3 and one-third

OpenStudy (anonymous):

explain it in word form thats what it means

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

\[6\frac{ 1 }{ 9 }-3\frac{ 1 }{ 3 }=?\]

OpenStudy (anonymous):

\[6\frac{ 1 }{ 9} - 3\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

\[55/9 - 10/3\]

OpenStudy (anonymous):

\[= 55/9 - 30/9\]

OpenStudy (anonymous):

= 25/9

OpenStudy (anonymous):

understood how I got here?

OpenStudy (anonymous):

can u tell ,e in words not in numberrs pl z

OpenStudy (anonymous):

tell me the answer in words tell me how i do it in words plz

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

6 1/9 is the same as 6 + 1/9 right?

OpenStudy (anonymous):

which is the same as 6/1 + 1/9

OpenStudy (anonymous):

just tell me the answer and how did u get it ok

OpenStudy (anonymous):

I got it by the method I used above

OpenStudy (anonymous):

i donht understand that just write it all in a paragragh not step ny step write it all in one sentence plz